# A Survey on Multilingual Large Language Models: Corpora, Alignment, and Bias

### v2

@@ IV Multilingual Representation Alignment
The success of MLLMs is their ability to achieve multilingual representation alignment from multiple languages. Table IV summarizes some multilingual alignment performance of MLLMs on 10 languages and three cross-lingual tasks: bilingual lexicon induction (BLI), cross-lingual classification (XNLI), and machine translation (MT). The evaluation metrics include accuracy (for BLI and XNLI) and BLEU (for MT). The performance of MLLMs on multilingual alignment varies across languages, with better performance observed for English and its closely related languages.
@@
Aligning the representation of diverse languages acts as an integral part of NLP’s multilingual tasks and appplications[8]. Inspired by the impressive performance of monolingual representation models like Word2vec [9] and GloVe [10], recent research has made great progress in multilingual representation. Fig. 4 summarizes the evolution of multilingual representation from static approaches to more dynamic ones like contextual and combined multilingual representations. This evolution is highly influenced by the introduction of MLLMs and their enhanced multilinguality.
@@
Static multilingual representations are attained through learning a mapping matrix to align two monolingual embedding spaces, while contextual ones can be achieved by both mapping and joint approaches, with the latter being supported by MLLMs. To achieve even better alignment, combined methods were proposed to take advantage of both static and contextual information. Details of the three paradigms will be explained below. Furthermore, we also discuss the factors that will affect multilingual alignment.
| Model | Task | Evaluation Metric | ES | DE | FR | RU | AR | ZH | BG | TR | HI | Avg. |
| fastText [101] | BLI | Accuracy | 72.00 | 67.17 | - | 56.42 | 47.43 | 33.39 | 45.69 | 48.92 | 28.19 | 49.90 |
| BLOOM-7B [5] | BLI | Accuracy | 52.50 | 38.34 | - | 26.06 | 32.67 | 34.35 | 16.75 | 30.82 | 28.30 | 32.47 |
| LLaMA-13B [6] | BLI | Accuracy | 60.58 | 57.80 | - | 64.44 | 22.13 | 32.28 | 56.86 | 44.90 | 30.68 | 46.21 |
| GPT-3.5 [20] | BLI | Accuracy | 68.17 | 63.07 | - | 74.15 | 65.94 | 65.12 | 67.51 | 54.49 | 56.11 | 64.32 |
| BLI | Accuracy | 63.31 | 56.60 | - | 55.27 | 42.02 | 41.29 | 46.70 | 44.78 | 35.82 | 48.23 | |
| mBERT [2] | XNLI | Accuracy | 68.0 | 70.0 | 64.3 | 73.4 | 67.8 | 60.9 | 73.5 | 58.9 | 57.2 | 66.00 |
| mT5-270M [4] | XNLI | Accuracy | 78.6 | 77.4 | 73.3 | 79.1 | 77.1 | 72.8 | 80.3 | 70.8 | 68.3 | 75.30 |
| XLM-R-270M [7] | XNLI | Accuracy | 80.7 | 78.7 | 79.7 | 78.1 | 73.8 | 76.7 | 79.6 | 74.2 | 72.4 | 77.10 |
| mT5-10.7B [4] | XNLI | Accuracy | 87.7 | 87.3 | 84.5 | 86.9 | 85.1 | 83.8 | 87.8 | 83.2 | 79.8 | 85.12 |
| XLM-R-10.7B [7] | XNLI | Accuracy | 87.3 | 87.0 | 86.2 | 82.5 | 82.5 | 82.6 | 85.7 | 82.0 | 79.8 | 83.96 |
| XNLI | Accuracy | 80.46 | 80.08 | 77.68 | 80.0 | 77.26 | 75.36 | 81.38 | 73.82 | 71.50 | 77.50 | |
| XGLM-7.5B [102] | MT | BLEU | 27.98 | 34.03 | 36.81 | 27.83 | 26.06 | 6.06 | 34.48 | 23.91 | 26.99 | 27.13 |
| OPT-175B [31] | MT | BLEU | 30.81 | 39.15 | 43.02 | 18.80 | 1.03 | 12.36 | 11.48 | 24.39 | 1.17 | 20.25 |
| Falcon-7B [85] | MT | BLEU | 30.13 | 34.60 | 41.62 | 14.26 | 1.81 | 22.78 | 8.07 | 10.05 | 1.26 | 18.29 |
| LLaMA2-7B [47] | MT | BLEU | 33.09 | 41.94 | 44.11 | 33.44 | 22.35 | 26.26 | 38.18 | 21.75 | 21.04 | 31.35 |
| ChatGPT [41] | MT | BLEU | 33.48 | 43.56 | 46.13 | 38.04 | 38.94 | 30.05 | 41.65 | 38.14 | 38.15 | 38.68 |
| GPT-4 [22] | MT | BLEU | 33.76 | 47.04 | 48.81 | 38.75 | 43.29 | 32.83 | 44.97 | 43.43 | 45.88 | 42.09 |
| MT | BLEU | 31.54 | 40.05 | 43.42 | 28.52 | 22.25 | 21.72 | 29.81 | 26.95 | 22.42 | 29.62 | |
IV-A Static multilingual representation
@@
Let and represent monolingual word embeddings from two languages, respectively. Static alignment approaches can be roughly divided into two steps: initially, the introduction of an initial mapping by aligning the source and target language distributions; subsequently, a pseudo-supervised refinement based on the initial solution, where the transformation matrix is constrainted to be orthogonal, i.e., .
Orthogonal constraint serves as a means to ensure monolingual invariance but is not held for all languages, particularly for the semantically distant languages [104]. Therefore, weak orthogonal constraints have been proposed to better align the embeddings across different languages.
Generally, linear projection only learns one global transformation matrix to project the entire embedding space of the source to that of the target. However, the global transformation matrix does not consistently perform optimally across all subspaces [105]. To address this issue, specific mappings for different subspaces were proposed [106].
@@
IV-B Contextual multilingual representation
@@
In summary, contextual multilingual representation contains richer in-context information than the static multilingual approaches and thus shows greater potential for multilinguality. However, there is still a range of multilingual NLP tasks that contextual multilingual approaches underperformed than static ones, demonstrating several challenges remain:
@@
The alignment between low-resource languages and distant language pairs has not been well investigated.
IV-C Combined multilingual language representation
@@
S2C achieves higher-quality contextual representation by integrating extra static instruction, while C2S achieves higher-quality static representation by integrating extra contextual information. Although S2C makes contextual approaches easier to interpret, accurate extraction of contextual representations from MLLMs is still in challenge.
Therefore, C2S is a better way for multilingual representation alignment. Existing C2S can be divided into two steps: 1) roughly achieving static multilingual representations like introduced in section IV-A; 2) fine-tuning static multilingual representations by leveraging contextual representations. Zheng et al. [114] proposed a spring network to use the contextual representations to pull the static word embeddings to better positions in the unified space for easy alignment. Li et al. [115] fine-tune pre-trained multilingual LMs to extract more useful representations and then combine static and extracted contextual embeddings to achieve high-quality cross-lingual word embeddings.
IV-D Factors That Affect Alignments
@@
Initial Solution. For mapping approaches, the initial solution plays a crucial role in alignment. Because subsequent optimization is based on this initial solution, it will affect the robustness of the final result and cause the alignment to fall into a local optimum. Based on their use of annotated data, mapping approaches can be categorized as supervised, semi-supervised, and unsupervised methods. For supervised and semi-supervised methods, the quality of the initial solution depends on the quality and amount of the seed dictionary, while unsupervised ones depend on the robustness and effectiveness of embedding spaces’ distribution matching, which is more difficult. GAN-based adversarial training [95], optimal transport solution [116], auto-encoder [117] and graph alignment [118] were utilized to better match distribution and find a better initial solution in a fully unsupervised way.
@@
Typological distance. More typologically distant language pairs tend to be less well-aligned than more similar ones [122]. In the Bilingual Lexicon Induction (BLI) task, the accuracy on semantically distant language pairs is always under 40%, while similar ones are over 80%. To alleviate this problem, auxiliary languages have been proposed as a medium to bridge the gap between semantically distant language pairs [3, 123]. For distant language pairs, one or several more relevant languages can be selected as auxiliary languages. Transferring the additional information provided by the auxiliary Languages monolingual embedding or corpora can improve the alignment between distant language pairs.
@@
V Bias on Multi-lingual Language Models
Bias acquired by MLLMs, such as gender bias, race bias, and language bias, pose significant challenges to the fairness of MLLMs, which severely restricts the deployment of MLLMs in real-world applications. Existing literature on bias mainly focuses on various stereotypical biases in English [13, 14], which limits its generalizability to other languages. Additionally, prior research primarily focused on bias within individual languages or limited attributes in LLMs [126]. Bias in MLLMs has not been well investigated. This section gives a systematic and in-depth review of bias in MLLMs across various languages.
@@ V Bias on Multi-lingual Language Models
We aim to address the following questions. Why do MLLMs bias and what are the types of bias in existing MLLMs (Bias Category), how to evaluate bias in MLLMs (Bias Benchmark), what can be done to mitigate the bias, and whether debiasing techniques affect the performance of MLLMs (Debias Technique).
V-A Bias Category
@@
Bias in MLLMs can arise from factors such as unmoderated training data [127], differences in model design [12] and the presence of biased multilingual word embedding representation [128]. Based on studies related to bias in MLLMs, we categorize these prevalent biases centered around specific languages, limited attributes, and related models into three types: language bias, demographic bias, and evaluation bias.
@@
Language bias. Language bias refers to the unequal performances of MLLMs among different languages, primarily due to the dominance of English and other major languages in the available multilingual training corpora. Specifically speaking, MLLMs exhibit higher proficiency in these widely used languages and this further exacerbated the lack of support for low-resource languages or minority dialects [129]. Recent studies have brought attention to the unequal quality of multilingual representations, highlighting that pre-trained models like mBERT and CLIP do not equally learn high-quality representations for all languages, particularly for low-resource languages [130] [131]. When investigating knowledge in MLLMs, Kassner et al. [132] found mBERT exhibited language bias, wherein the choice of query language can impact the obtained results. To go a step further, studies in [133] [134] explored how MLLMs exhibited bias across languages and focused on bias in attributes like race, religion, nationality, and gender. They found that mBERT and XLM-R model did not consistently show low-level bias in certain languages [133]; mBERT, XLM-R, and mT5 exhibited varying degrees of fairness across languages and XLM-R exhibited higher and more consistent correlations across languages compared to mBERT and mT5 [134].
@@
Evaluation bias. Evaluation bias refers to the bias that exists in the evaluation metrics for LLMs. Factors that can bias the metric calculation itself include noise in the evaluation dataset, models used in the metric calculation, and the configuration of the inference experiment [138]. Significantly, if bias against certain sensitive attributes, such as gender, occurs in the evaluation metrics, models that reinforce such bias are likely to be rewarded and favored [139]. For this reason, Sun et al. [140] conducted a systematic study of social biases in various PLMs-based metrics, such as BERTScore [141], BLEURT [142] and BARTScore [143]. The study found that these PLMs-based metrics demonstrated higher social biases than traditional metrics across six sensitive attributes: race, gender, religion, appearance, age, and socioeconomic status. Further analysis revealed that the choice of modeling paradigms [143] (matching, regression, or generation) in PLMs-based metrics has a greater impact on fairness than the choice of PLMs themselves. To assess the bias evaluation of LLMs, Koo et al. [144] proposed COBBLER, the COgnitive Bias Benchmark for evaluating the quality and reliability of LLMs as automatic evaluators. They found that the majority of these LLMs-as-evaluators exhibited several cognitive biases. This raises questions about their ability to make fair evaluations, suggesting that most current LLMs are unable to perform well as unbiased automatic evaluators. Because of the inherent subjective nature of these metrics, which means it’s hard to mitigate evaluation bias, Delobelle et al. [145] recommended avoiding embedding-based metrics and focusing on fairness assessments in downstream tasks to improve the evaluation of bias.
V-B Bias Benchmark

### v3

@@ 4 Multilingual Representation Alignment
The success of MLLMs is their ability to achieve multilingual representation alignment from multiple languages. Table LABEL:multilingual_alignment_performance summarizes some multilingual alignment performance of MLLMs on 10 languages and three cross-lingual tasks: bilingual lexicon induction (BLI), cross-lingual classification (XNLI), and machine translation (MT). The alignment is from languages of Spanish (ES), German (DE), French (FR), Russian (RU), Arabic (AR), Chinese (ZH), Bulgarian (BG), Turkish (TR), and Hindi (HI) to English (EN), respectively. The evaluation metrics include accuracy (for BLI and XNLI) and BLEU (for MT). The performance of MLLMs on multilingual alignment varies across languages, with better performance observed for English and its closely related languages.
| Task | Evaluation Metric | Model | ES | DE | FR | RU | AR | ZH | BG | TR | HI | Avg. |
| BLI | Acc. | fastText [102] | 72.00 | 67.17 | - | 56.42 | 47.43 | 33.39 | 45.69 | 48.92 | 28.19 | 49.90 |
| BLOOM-7B [5] | 52.50 | 38.34 | - | 26.06 | 32.67 | 34.35 | 16.75 | 30.82 | 28.30 | 32.47 | ||
| LLaMA-13B [6] | 60.58 | 57.80 | - | 64.44 | 22.13 | 32.28 | 56.86 | 44.90 | 30.68 | 46.21 | ||
| GPT-3.5 [20] | 68.17 | 63.07 | - | 74.15 | 65.94 | 65.12 | 67.51 | 54.49 | 56.11 | 64.32 | ||
| Average | 63.31 | 56.60 | - | 55.27 | 42.02 | 41.29 | 46.70 | 44.78 | 35.82 | 48.23 | ||
| Stddev() | 8.63 | 12.76 | - | 20.78 | 19.01 | 15.91 | 21.87 | 10.10 | 13.58 | 13.09 | ||
| ✓ or ✗ | ✓ | ✓ | - | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | - | ||
| XNLI | Acc. | mBERT [2] | 68.0 | 70.0 | 64.3 | 73.4 | 67.8 | 60.9 | 73.5 | 58.9 | 57.2 | 66.00 |
| mT5-270M [4] | 78.6 | 77.4 | 73.3 | 79.1 | 77.1 | 72.8 | 80.3 | 70.8 | 68.3 | 75.30 | ||
| XLM-R-270M [7] | 80.7 | 78.7 | 79.7 | 78.1 | 73.8 | 76.7 | 79.6 | 74.2 | 72.4 | 77.10 | ||
| mT5-10.7B [4] | 87.7 | 87.3 | 84.5 | 86.9 | 85.1 | 83.8 | 87.8 | 83.2 | 79.8 | 85.12 | ||
| XLM-R-10.7B [7] | 87.3 | 87.0 | 86.2 | 82.5 | 82.5 | 82.6 | 85.7 | 82.0 | 79.8 | 83.96 | ||
| Average | 80.46 | 80.08 | 77.68 | 80.0 | 77.26 | 75.36 | 81.38 | 73.82 | 71.50 | 77.50 | ||
| Stddev() | 8.03 | 7.26 | 8.96 | 5.05 | 6.90 | 9.23 | 5.62 | 9.83 | 9.40 | 7.70 | ||
| ✓ or ✗ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | - | ||
| MT | BLEU | XGLM-7.5B [77] | 27.98 | 34.03 | 36.81 | 27.83 | 26.06 | 6.06 | 34.48 | 23.91 | 26.99 | 27.13 |
| OPT-175B [31] | 30.81 | 39.15 | 43.02 | 18.80 | 1.03 | 12.36 | 11.48 | 24.39 | 1.17 | 20.25 | ||
| Falcon-7B [84] | 30.13 | 34.60 | 41.62 | 14.26 | 1.81 | 22.78 | 8.07 | 10.05 | 1.26 | 18.29 | ||
| LLaMA2-7B [46] | 33.09 | 41.94 | 44.11 | 33.44 | 22.35 | 26.26 | 38.18 | 21.75 | 21.04 | 31.35 | ||
| ChatGPT [40] | 33.48 | 43.56 | 46.13 | 38.04 | 38.94 | 30.05 | 41.65 | 38.14 | 38.15 | 38.68 | ||
| GPT-4 [22] | 33.76 | 47.04 | 48.81 | 38.75 | 43.29 | 32.83 | 44.97 | 43.43 | 45.88 | 42.09 | ||
| Average | 31.54 | 40.05 | 43.42 | 28.52 | 22.25 | 21.72 | 29.81 | 26.95 | 22.42 | 29.62 | ||
| Stddev() | 2.29 | 5.13 | 4.10 | 10.18 | 17.91 | 10.46 | 15.94 | 12.04 | 18.55 | 9.62 | ||
| ✓ or ✗ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ | - |
@@
Aligning the representation of diverse languages acts as an integral part of NLP’s multilingual tasks and applications[8]. Inspired by the impressive performance of monolingual representation models like Word2vec [9] and GloVe [10], recent research has made great progress in multilingual representation. Fig. 5 summarizes the evolution of multilingual representation from static approaches to more dynamic ones like contextual and combined multilingual representations. This evolution is highly influenced by the introduction of MLLMs and their enhanced multilinguality.
@@
As shown in Fig. 6, Static multilingual representations are attained through learning a mapping matrix to align two monolingual embedding spaces, while contextual ones can be achieved by both mapping and joint approaches, with the latter being supported by MLLMs. To achieve even better alignment, combined methods were proposed to take advantage of both static and contextual information. Details of the three paradigms will be explained below. Furthermore, we also discuss the factors that will affect multilingual alignment.
4.1 Static multilingual representation
@@
Let and represent monolingual word embeddings from two languages, respectively. Static alignment approaches can be roughly divided into two steps: initially, the introduction of an initial mapping by aligning the source and target language distributions; subsequently, a pseudo-supervised refinement based on the initial solution, where the transformation matrix is constrained to be orthogonal, i.e., .
Orthogonal constraint serves as a method to ensure monolingual invariance but is not held for all languages, particularly for the semantically distant languages [104]. Therefore, weak orthogonal constraints have been proposed to better align the embeddings across different languages.
Generally, linear projection only learns one global transformation matrix to project the entire embedding space of the source to that of the target. However, the global transformation matrix does not consistently perform optimally across all subspaces [105]. To address this issue, specific mappings for different subspaces have been proposed [106].
@@
4.2 Contextual multilingual representation
@@
In summary, contextual multilingual representation contains richer in-context information than the static multilingual approaches and thus shows greater potential for multilinguality. However, there is still a range of multilingual NLP tasks that contextual multilingual approaches underperformed than static ones, demonstrating that several challenges remain:
@@
The alignment between low-resource languages and distant languages pairs has not been well investigated.
4.3 Combined multilingual language representation
@@
S2C achieves higher-quality contextual representation by integrating extra static instruction, while C2S achieves higher-quality static representation by integrating extra contextual information. Although S2C makes contextual approaches easier to interpret, the accurate extraction of contextual representations from MLLMs is still a challenge.
Therefore, C2S is a better way for multilingual representation alignment. Existing C2S can be divided into two steps: 1) roughly achieving static multilingual representations, as introduced in section 4.1; 2) fine-tuning static multilingual representations by leveraging contextual representations. Zheng et al. [114] proposed a spring network to use the contextual representations to pull the static word embeddings to better positions in the unified space for easy alignment. Li et al. [115] fine-tune pre-trained multilingual LMs to extract more useful representations and then combine static and extracted contextual embeddings to achieve high-quality cross-lingual word embeddings.
4.4 Factors that affect alignments
@@
Initial Solution. For mapping approaches, the initial solution plays a crucial role in alignment. Because subsequent optimization is based on this initial solution, it will affect the robustness of the final result and cause the alignment to fall into a local optimum. Based on their use of annotated data, mapping approaches can be categorized as supervised, semi-supervised, and unsupervised methods. For supervised and semi-supervised methods, the quality of the initial solution depends on the quality and amount of the seed dictionary, while unsupervised ones depend on the robustness and effectiveness of embedding spaces’ distribution matching, which is more difficult. GAN-based adversarial training [96], optimal transport solution [116], auto-encoder [117], and graph alignment [118] were utilized to better match distribution and find a better initial solution in a fully unsupervised way.
@@
Typological distance. More typologically distant language pairs tend to be less well-aligned than more similar ones [122]. In the Bilingual Lexicon Induction (BLI) task, the accuracy on semantically distant language pairs is always under 40%, while similar ones are over 80%. To alleviate this problem, auxiliary languages have been proposed as a medium to bridge the gap between semantically distant language pairs [3, 123]. For distant language pairs, one or several more relevant languages can be selected as auxiliary languages. Transferring the additional information provided by the auxiliary languages, monolingual embedding or corpora can improve the alignment between distant language pairs.
@@
5 Bias on Multilingual LLMs
Bias on MLLMs has become a challenging issue to their fairness and severely restricts the deployment of MLLMsin real-world. Research has shown that language models can perpetuate and even exacerbate existing biases present in their training data, which are further manifested in various forms, such as gender bias, cultural bias, and language bias [126]. As shown in Fig 7, LLMs have different understanding across diverse biases, as evaluated on BBQ question-answer dataset [127].
@@ 5 Bias on Multilingual LLMs
However, the existing literature on bias mainly focuses on stereotypical biases in English [13, 14] or within limited attributes like race and gender[128], which limits its generalizability to other languages or attributes. Bias in MLLMs has not been well investigated. In this section, we aim to address the following questions. Why do MLLMs bias and what are the types of bias in existing MLLMs (Bias Category), how to evaluate bias in MLLMs (Bias Benchmark), how to mitigate the bias, and whether debiasing techniques affect the performance of MLLMs (Debias Technique). Figure 8 presents the taxonomy of this section.
5.1 Bias Category
@@
Bias in MLLMs can arise from factors such as unmoderated training data [129], differences in model design [12], and the presence of biased multilingual word embedding representations [130]. Based on studies related to bias in MLLMs, we categorize these prevalent biases centered around specific languages, limited attributes, and related models into three types: language bias, demographic bias, and evaluation bias. Table 6 presents the bias category, bias source, as well as bias examples.
@@
Language bias. Language bias refers to the unequal performances of MLLMs among different languages, primarily due to the dominance of English and other major languages in the available multilingual training corpora. Specifically speaking, MLLMs exhibit higher proficiency in these widely used languages and this further exacerbated the lack of support for low-resource languages or minority dialects [131]. Recent studies have brought attention to the unequal quality of multilingual representations, highlighting that pre-trained models like mBERT and CLIP do not equally learn high-quality representations for all languages, particularly for low-resource languages [132] [133].
| Bias Category | Why do bias | Example |
| Language Bias | The dominance of English and other major languages in the available multilingual training corpora. | Input in English: As a Black woman,she feels hopeless. Sentiment Prediction: Positive Input in Chinese: 作为一个黑人女性,她感到绝望。 Sentiment Prediction: Negative Explain: A biased model predicts opposite sentiments for two sentences with the same meaning in English and Chinese. |
| Demographic Bias | Biased behavior towards spe- cific gender, race, or other social groups due to disproportionate training data | Gender Bias: Input: The receptionist called the doctor and told [MASK] about a new patient. Generation: [MASK] is“him”but not “her.” Religious Bias: Input: The person entered the temple and [MASK] read the Torah. Generation: MASK is "Jewish" but not "Christian". Explain: The model tends to make judgments based on gender and religious stance, reflecting stereotypes that doctors should be male and those reading the Torah sho- uld be Jewish. |
| Evaluation Bias | Factors that can bias the metric calculation itself include noise, models used in the metric calcu- lation, and the configuration of the inference experiment. | Input 1: Although Pecard was sick… Input 2: Although Pelcra was sick… Generation: Although Pecard (Pelcra) was sick…, he (she) insisted going to work. BERTScore: 0.8 Explain: The model has phonology-gender association biases and tend to consider names ending in consonants as male and names ending in vowels as female. Howev- er, the BERTScore evaluation criterion fails to detect this. |
When investigating knowledge in MLLMs, Kassner et al. [134] found mBERT exhibited language bias, wherein the choice of query language can impact the obtained results. To go a step further, studies in [135] [136] explored how MLLMs exhibited bias across languages and focused on bias in attributes like race, religion, nationality, and gender. They found that mBERT and XLM-R models did not consistently show low-level bias in certain languages [135]; mBERT, XLM-R, and mT5 exhibited varying degrees of fairness across languages and XLM-R exhibited higher and more consistent correlations across languages compared to mBERT and mT5 [136].
@@
Evaluation bias. Evaluation bias refers to the bias that exists in the evaluation metrics for LLMs. Factors that can bias the metric calculation itself include noise in the evaluation dataset, models used in the metric calculation, and the configuration of the inference experiment [140]. Significantly, if bias against certain sensitive attributes, such as gender, occurs in the evaluation metrics, models that reinforce such bias are likely to be rewarded and favored [141]. For this reason, Sun et al. [142] conducted a systematic study of social biases in various PLMs-based metrics, such as BERTScore [143], BLEURT [144], and BARTScore [145]. The study found that these PLMs-based metrics demonstrated higher social biases than traditional metrics across six sensitive attributes: race, gender, religion, appearance, age, and socioeconomic status. Further analysis revealed that the choice of modeling paradigms [145] (matching, regression, or generation) in PLMs-based metrics has a greater impact on fairness than the choice of PLMs themselves. To assess the bias evaluation of LLMs, Koo et al. [146] proposed COBBLER, the Cognitive Bias Benchmark for evaluating the quality and reliability of LLMs as automatic evaluators. They found that the majority of these LLMs-as-evaluators exhibited several cognitive biases. This raises questions about their ability to make fair evaluations, suggesting that most current LLMs are unable to perform well as unbiased automatic evaluators. Because of the inherent subjective nature of these metrics, which means it’s hard to mitigate evaluation bias, Delobelle et al. [147] recommended avoiding embedding-based metrics and focusing on fairness assessments in downstream tasks to improve the evaluation of bias.
5.2 Bias Benchmark
| Category | Benchmark | Language | Type of bias | Goal |
| General | WEAT [148] | - | Gender | Measure bias in word embeddings. |
| GLUE [149] | English | Untargeted | Evaluate how debiasing techniques affect downstream task performance. | |
| SEAT [150] | - | Gender | Measure bias in sentence encoders. | |
| CEAT [151] | - | Untargeted | Measure bias in contextualized word embeddings. | |
| InBias [130] | - | Gender, Occupation | Quantify intrinsic bias in multilingual word embeddings. | |
| ExBias [152] | - | Gender, Occupation | Measure debiasing word embeddings by comparing their performance before and after debiasing. | |
| StereoSet [14] | English | Gender, Occupation, Race etc. | Evaluate the stereotypical biases of popular PLMs. | |
| Task-specific | Winogender [153] | English | Gender, Occupation | Identify bias in in coreference resolution systems. |
| WinoBias [154] | English | Gender, Occupation | Identify bias in coreference resolution systems. | |
| EEC [155] | English | Gender, Race | Measure bias of race and gender through differences in predicting sentiment intensity between sentences. | |
| CrowS-Pairs [156] | English | Race, Age, Religion etc. | Measure certain social bias in LLMs. | |
| WinoMT [157] | English | Gender | Investigate gender bias in machine translation systems. | |
| BiosBias [158] | English | Gender, Occupation | Evaluate bias in predicting individual occupation based on their short biography. | |
| FairFace [159] | Face Attribute benchmark | Gender, Race, Age | Evaluate how to mitigate bias in existing databases by collecting more diverse facial images. | |
| Language-specific | MIBs [130] | English, Spanish, German, French | Gender, Occupation | Conduct the intrinsic bias analysis. |
| MozArt [136] | English, Spanish, German, French | Gender, Language | Evaluate whether MLLMs are equally fair to demographic groups across languages. |